### v4

@@ Abstract
Intrusion Detection Systems (IDS) are widely employed to detect and mitigate external network security events. Vehicle ad-hoc Networks (VANETs) are evolving, especially with the development of Connected Autonomous Vehicles (CAVs). It is, therefore, crucial to assess how traditional IDS approaches can be utilised for these emerging technologies. We combat this issue by introducing a stacked ensemble learning scheme for IDS to improve the detection of threats. Our method employs several machine learning (ML) algorithms to identify threats more precisely than single methods. We demonstrate our approach’s performance compared to current ML methods on CICIDS2017 and VeReMi and establish that it can better discriminate threats. Our process also incorporates hyperparameter optimisation and feature selection to improve performance further. Overall, our results suggest that stacked ensemble learning is a promising technique for enhancing the effectiveness of IDS.
@@ 1 Introduction
Vehicles are increasingly being connected to their operating environments, resulting in Vehicle Ad-hoc Networks being considered a subset of Mobile Ad-hoc Networks when conditions are altered and environmental changes occur due to the continual rotation of automobiles. Inter-vehicle communication is a growing research, standardisation, and development area as more cars are configured with wireless communication and computing technology. The new era of automobiles and road traffic systems will depend on wireless networks and the associated data to better coordinate and manage traffic flow safely and effectively through developments such as machine learning models. Ultimately, the innovations of VANETs must enable safer and more efficient travel for road users. To enable this innovation, we need a mechanism by which vehicles can be made sufficiently intelligent to manage their process for road safety. This notion constituted the foundation for VANETs. Software-defined networking (SDN) based VANETs follow the below architecture Ku et al. (2014). VANET is a wireless network in which both stationary and mobile nodes connect. Such a system has frequent disconnections, a dynamic topology, and significant node mobility. Each attribute may result in a security flaw that hazards the system. To detect known and unencrypted attacks in a local detection scenario, researchers are showing great interest in current years due to the SDN concept in VANET that implements the ML-based detection mechanism. IDS is still in its infancy, and many studies must be conducted to make ID as successful as possible. Current IDS is plagued by difficulties and obstacles that require rapid and intensive investigation. This work has identified significant elements and activities that research communities must address. The problem and assignment areas :
  • Incomplete data set: A data set is a collection of comprehensive data and information from a survey that must be examined. From IDSs, data sets contribute significantly to the accuracy of outcomes. For example, researchers are now utilising DARPA 98, 99, and the University of New Mexico’s immune system data sets. Still, we cannot prevent relatively recent attacks due to their lifetime.
  • Methodologies for detection: It is a crucial aspect of determining if a packet/data originates from an attack or whether it is valuable data that the user needs for carrying out operations or activities. Detection algorithms must be efficient enough to match all cases in a short amount of time and efficiently match criteria. The basis for the search policy might be abuse or abnormality. Therefore, anomaly or abuse might serve as the detection approach. A declaration is a piece of data utilised by the matching process. A downside is that there are no instructions for matching fresh assaults. Therefore, this attack is detected by determining if a new attack has been discovered or by detecting a change in data that matches the pattern. Thus, we need a good and quick algorithm that will identify patterns wider and faster to fit the majority of assaults.
  • Weak design: All IDSs have a compact design; thus, if a user needs to edit a portion of the IDS, the IDS must be stopped, modified, and redeployed.
@@
By referring to the literature review below, several essential study gaps in ML-based IDS research have been noted. Still, there is enough opportunity to improve the accuracy of the intrusion detection model. Moreover, IDS users still struggle to detect known unencrypted attacks because IDS has not yet achieved 100% accuracy. One significant issue with IDS is that they regularly alert the auditor to false positives. However, an IDS can be optimised to reduce the number of false positives and false negatives, which is a deficiency of publicly available VANET databases with attacks. No fast, inexpensive, and effective in-built privacy-preserving ML-based IDS is still visible.
@@ 1.1 Contributions
This article proposes a new model for anomaly detection in the SDN-based VANET for the Binary and Multiclass classification to detect different attacks and Benign in the VeReMi dataset. The contributions of this paper are as follows:
@@
    This research suggests a novel approach to intrusion detection in SDN-based VANETs that integrates different machine-learning approaches for better accuracy and speed of operation.
@@
    Furthermore, the suggested IDS model can identify minor attack entries in a big dataset, an essential component of VANET security.
  3. 3.
    The evaluation of our suggested approach using two comprehensive datasets (CICIDS2017 and VeReMi) can provide a more comprehensive evaluation of IDS models in SDN-based VANETs. VeReMi offers a large and diverse network traffic data set, while CICIDS2017 focuses on intrusion detection in a simulated environment relevant to VANET security. This combination can help evaluate IDS performance in detecting and responding to a broader range of security threats. Insights into the strengths and weaknesses of the proposed ML model and techniques can also be gained.
@@ 2 Background
VANET, a type of wireless network connecting both stationary and mobile nodes, is characterised by frequent disconnections, a dynamic topology, and high mobility, which can lead to potential security flaws. Compared to Mobile Ad Hoc Networks (MANET), VANETs exhibit considerably greater mobility and more stringent topology restrictions. Vehicular Ad Hoc Networks (VANETs) enable applications such as real-time traffic monitoring, blind crossing prevention, dynamic route planning, collision prevention, and internet access for vehicles. According to the World Health Organization (WHO), 1.3 million people die in road crashes annually, with 20 to 50 million suffering non-fatal injuries WHO (2023).
Figure 1: Software-defined VANET
The new era of automobiles and road traffic systems will depend on wireless networks and machine learning. As stated previously, many lives have been lost on highways. Consequently, when road traffic increases, road safety must also grow. We need a mechanism by which vehicles can be made sufficiently intelligent to manage their process for road safety. This notion constituted the foundation for VANETs. SDN-based VANETs follow the below architecture Ku et al. (2014).
2.1 Communication And Standerd
The USDOT (The United States Department of Transportation) has designated Dedicated Short-Range Communication as the preferred inter-vehicle communication medium. The communication infrastructure for linked automobiles is provided via WAVE protocols. IEEE 802.11p protocol and IEEE 1609.x protocols define WAVE. Their activities conform to the internet protocol stack. Table I displays WAVE’s tiered architecture. Physical and data connection layers are supported by IEEE 802.11p. The tiers of operation for IEEE 1609.x are network, transport, and application. This protocol focuses on the physical and Media Access Control (MAC) levels to provide low latency and good dependability across short radio communication lines. Chekkouri et al. (2015) It functions at both the physical and MAC levels. It offers a direct wireless interface and mapping between MAC and physical data. In addition, it provides media access rules and aids in multi-channel operations at the MAC layer.
@@ 2.1 Communication And Standerd
enables security via authentication and encryption of private communications. All cars are expected to be registered by a Motor Vehicle registration unit. This would assign MAC IDs to automobiles. Alternatively, Vehicle Identification Numbers (VINs) can serve as MAC addresses for automobiles. Vehicles will obtain IP addresses upon network connection. Vehicle identification will be based on their MAC addresses, IP addresses, or publicly provided IDs. (Zheng and Wu, 2016)
@@
Support for multi-channel operations is achieved by time division so radios can switch between different channel intervals. Control Channel (CCH) and Service Channel (SCH) are the two channels of this IEEE standard. Continuous safety messages are sent via CCH, whilst other applications are routed via SCH. IEEE 1609.3p defines the network and transport layer services. Here, IP addressing and routing occur. IEEE 1609.2p is the protocol that specifies secure message exchanges and encryption. IEEE 1609.1p is a protocol specifying alert and probe vehicle message forms. Currently, a modification to IEEE 1609.x is in the works. The following enhancements have been made. Draft IEEE 1609.5 to handle vehicle-to-vehicle and vehicle-to-infrastructure connectivity. Draft IEEE 1609.6 to provide remote management between Onboard Units and Road-Side-Units. IEEE Standard 1609.11 for safe electronic transactions.IEEE Standard 1609.12 defines the allocation of value identifiers in WAVE. (Chen, Jiang and Delgrossi, 2009)
2.2 SDN strategies to address VANET security concerns
@@
2.3 VANET Characteristics and Negatives
VANET is a wireless network in which both stationary and mobile nodes connect. Such a system has frequent disconnections, a dynamic topology, and significant node mobility. Each attribute may result in a security flaw that hazards the system. Due to the high mobility of nodes, VANET topology is dynamic and unpredictable; this and other circumstances, such as severe weather conditions, may cause VANET to have several disconnections throughout the system.
@@ 3 Related Works
As a part of the research, different articles have been reviewed to explain how earlier researchers have ended work on an Intrusion Detection System for cyber-attack detection in SDN-based VANET. This Literature review study employs the methods stated by Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA). The methodology covers eight main steps: User needs assessment, define the Research Questions, Search String, Selection of Source, Select Relevant Paper, Screening Findings, Extract Data, and finally, synthesise the collected data. (Kitchenham and Charters, 2007)
3.1 User Needs Assessment
To effectively determine user requirements for cyber-attack protection, the following essential steps should be taken:
3.1.1 Step 1: Identify Stakeholders
Government and private organisations are the two main stakeholders in the cybersecurity field. In the private sector, companies from various sectors of the economy, the ISPs, software vendors, end users etc.
Table 1: CICIDS2017 Dataset attacks.
Attacks Size
DoS 251712
PortScan 158804
DDoS 128025
Brute Force 13832
Web Attack 2180
Bot 1956
Infiltration 36
Heartbleed 11
@@ 3.1.1 Step 1: Identify Stakeholders
3.1.2 Step 2: Understand IDS Issues
IDS and Malware detection systems can still not detect unknown, Encrypted, and Inference attacks with 100% accuracy. According to Kaspersky, most Malware detection products do not even reach the required 90% security threshold. A significant player in the IDS market, CISCO, claimed that Rapid detection—those made in less than five minutes—and long-term learning capacities displayed excellent performance in their IDS. Cisco Encrypted Traffic Analytics, or IDS, discovered nearly two-thirds of all known malicious flows in less than five minutes. Moreover, Cisco declared this IDS demonstrated better identification results even with low volumes of 0 to 20 flows, whereas 2000 plus flows enabled 100% detection for known attacks. Nevertheless, this solution has a limitation with a small attack size and taking three hours to gain 100% accuracy. However, the Cisco Adaptive Security Device Manager (ASDM) to manage Cisco Secure Firewall ASA (with IDS/IPS) has been impacted by Common Vulnerabilities and Exposures CVE-2022-20829, CVE-2021-1585, and CVE-2022-20828. Researcher Jake Baines, who disclosed them to Cisco in February and March of 2022, added that many users were undoubtedly not updating their Cisco firewalls properly. CISCO (2022) Scroxton (2022).
3.2.1 CICIDS2017
The CICIDS2017 dataset offers a benchmark for evaluating intrusion detection systems with various network characteristics and attack scenarios, including Brute Force, Heartbleed, Botnet, DoS, DDoS, Web Attack, and infiltration. The dataset collected over five days in July includes over 80 traffic characteristics per flow, obtained using the CICFlowMeter tool from the Canadian Institute for Cybersecurity. However, this dataset may not cover all authentic network behaviours due to individual differences in attackers’ strategies and timings. Sharafaldin et al. (2018). Nevertheless, CICIDS2017 holds more than 280,000 entries with all unfilled features. Very few numbers of entries contain ’NaN’ or ’Infinity.’
3.2.2 VeReMi
The VeReMi dataset is a synthetic benchmark dataset for misbehaviour detection systems in Vehicular Ad Hoc Networks that relies on the VEINS framework addendum. VeReMi is implemented with the VEINS framework, modified from Version 4.6, with the help of LuST (Luxembourg Simulation of Urban Mobility Traffic). This dataset can be efficiently employed to test the performance of misbehaviour detection approaches implemented on a city-wide scale while utilising fewer computing resources than standard (frequent) VEINS simulations. The dataset comprises On-Board Unit (OBU) message logs, including timestamps, sender IDs, positions, speeds, and reception times. VeReMi also includes a verification file, which provides GUI parameters on the validity of the data of each message upon which the verification confirms whether it is a "normal" or "attack" case. van der Heijden et al. (2018)
The dataset spreads across three attacker densities, five distinctive attack types, and three levels of density; it also includes various scenarios at different times of the day, including how traffic varies across the time of the day with the vehicle density range from 35 vehicles in low-dense areas to 519 vehicles during maximum congestion; and diverse types of attacks where the attacks in this paper are attacks against the position expressed in many ways from the position remaining still to noising out the position. While versatile, the VeReMi dataset is restricted in terms of both position-falsifying attacks and all types of network threats and advanced malicious strategies. Being a synthesised dataset, it lacks real-life cases but can be developed further by researchers.
The dataset is publicly available for reproducibility; thus, the research results made from the dataset are reliable and verifiable. VeReMi has 562,467 entries, out of which 437,429 entries are benign and 125,038 are attacking entries. The dataset comprehensively covers a city-level context of various types of roads and traffic behaviours, thus making it a reliable framework for benchmarking and testing the performance of misbehaviour detection approaches in VANETs. van der Heijden (2018)
@@ 3.2.3 Attack detection techniques
The article discusses the ML-based tactic to detect DDoS Attacks in SDVN by Singh et al. (2018). This paper suggests a method for spotting threats in SDN-based vehicular networks based on machine learning. The research has a particular emphasis on vehicle-to-infrastructure (V2I) links. It analyses the current environment to establish the ML strategy that will most successfully detect threats posed by distributed denial of service (DDoS) attacks (Short-lived Transmission Control Protocol Flooding, User Datagram Protocol Flooding). To accomplish this, it collects seven different elements of actual data from TCP (Transmission Control Protocol) and UDP (User Datagram Protocol) flows both when DDoS attacks them and when they are not. The network is educated by employing several different supervised learning algorithms depending on the data acquired, and it is discovered that the gradient boost classifier produces the most satisfactory outcomes. In addition, the random forest, decision tree, and linear support vector machine (linear-SVM) approach all produce results that are close to being the best. On the other hand, the limited training data could be to blame for the weak performance of Neural Networks (NN). Accurate, measurable data was not found in the write-up or the chart.
@@
Aneja et al. (2018) presents an AI-based system to identify a Route Request (RREQ) flooding attack. Researchers used a genetic algorithm (GA) to optimise feature subset selection and an Artificial Neural Network (ANN) for classification. The model scored 99% accuracy from Network Simulator 2 (NS2) 1000 entries simulated dataset.
@@
Yu et al. (2018) In this study, they analyse DDoS assaults in an SDN-based vehicular network by employing TCP flood, UDP flood, or ICMP (internet control message protocol) flood. In this research paper, the authors present a scheme for detecting Distributed Denial-of-Service (DDoS) attacks in SDVN atmospheres using three separate models. A "trigger detection model" for inbound packets and another method is the "flow table feature-based detection model." Both methods use the OpenFlow protocol features and the detection of a threat model based on support vector machine classification. All of these models are used in conjunction with one another. They synthesised and simulated network traffic to accommodate various DDoS attacks via Scapy and hping3, and the findings suggested an efficiency of over 97%. In the article, ML-Based Intrusion Detection Method for VANETs.
@@
Zeng et al. (2018) the authors developed an ML-based Senior2Local IDS with propagation to identify potential threats in VANETs both Globally and Locally. They used a fuzzification ANN technique to recognise malicious attacks globally and protect Cluster Heads in Roadside Units (RSUs). After that, a game theory trust system or RSU is developed separately. They employed a simple and straightforward Support Vector Machine (SVM) to detect attacks locally. The findings demonstrated that the method was more reliable and resilient than earlier ML-based tactics.
@@
Gad et al. (2021) conduct the experiment on a ToN-IoT dataset and test various ML approaches in binary and multiclass classification problems in VANET to detect intrusion. On the other hand, LR is used for solving binary classification problems, while multiclass issues are addressed with a one-vs-rest approach.
@@
3.2.4 Each method has yet to reach optimum accuracy
Türkoğlu et al. (2022), With the help of feature selection and hyperparameter optimisation, the DT classifier attained an accuracy score of 99.35%, which was the best possible result. The excellent results achieved with the methodology developed proved the considerable significance of feature selection and optimisation techniques. These findings were demonstrated by the outstanding results obtained with the method.
Kim et al. (2017) proves the existence of a collaborative security threat detection method within software-defined vehicular cloud architecture. A traditional vehicular environment has some exceptional features, such as high flexibility, dynamic networking, and rapid connectivity. His proposed model said a Speed Dependent Volume Controller (SDVC) had collected the flow information from the vehicular cloud (VC) to train multiclass SVM. In the next step, the SVM classifier identifies the bad actors among the network’s nodes. However, performance is still under 90% for Accuracy score, Precision, and Recall.

### v5

@@ Abstract
Intrusion Detection Systems (IDS) are widely employed to detect and mitigate external network security events. Vehicle ad-hoc Networks (VANETs) continue to evolve, especially with developments related to Connected Autonomous Vehicles (CAVs). In this study, we explore the detection of cyber threats in vehicle networks through ensemble-based machine learning, to strengthen the performance of the learnt model compared to relying on a single model. We propose a model that uses Random Forest and CatBoost as our main ‘investigators’, with Logistic Regression used to then reason on their outputs to make a final decision. To further aid analysis, we use SHAP (SHapley Additive exPlanations) analysis to examine feature importance towards the final decision stage. We use the Vehicular Reference Misbehavior (VeReMi) dataset for our experimentation and observe that our approach improves classification accuracy, and results in fewer misclassifications compared to previous works. Overall, this layered approach to decision-making — combining teamwork among models with an explainable view of why they act as they do — can help to achieve a more reliable and easy-to-understand cyber security solution for smart transportation networks.
@@ 1 Introduction
As vehicles increasingly connect to their surroundings, Software-Defined Vehicle Ad-hoc Networks (SD-VANETs) are emerging as a vital subset of Mobile Ad-hoc Networks. These networks become particularly relevant in dynamic conditions where the environment shifts due to the constant movement of vehicles. The surge in inter-vehicle communication represents a crucial area of research, standardization, and development, propelled by the growing incorporation of wireless communication and computing technologies in automobiles. This integration forms the backbone of SD-VANETs (Figure 1), which facilitate wireless connections between both stationary and mobile nodes, enabling seamless communication that enhances road safety and efficiency.
Figure 1: Software-defined VANET
However, this advancement has a serious downside: it opens up a new battleground for cyber threats. Automotive cybersecurity is not just a technical challenge—it is a matter of public safety, privacy, and the stability of entire transportation systems. While traditional Intrusion Detection Systems (IDS) can catch many threats, they often fall short where it matters most: explaining why certain threats are flagged or missed, especially in fast-paced, real-time environments.
Interpretability in IDS models is essential. First, VANETs are dynamic, with high-speed vehicles and constantly shifting network conditions. Security systems for these environments need to be more than accurate—they need to be transparent, so operators can trust and quickly act on what the system detects. Second, a model that can explain itself allows designers to see exactly where it might go wrong, giving them a way to strengthen it and cut down on false positives. In high-traffic vehicle networks, even minor errors can lead to major problems, so interpretability is non-negotiable.
This study tackles this critical gap by introducing an innovative IDS framework that combines Explainable Artificial Intelligence (XAI) with SHAP (SHapley Additive exPlanations) analysis. This approach goes beyond simply detecting threats; it shows why certain patterns are detected, providing clarity and reliability in the complex landscape of VANET security. Our focus on interpretability alongside performance sets a new standard in cybersecurity for intelligent transportation, meeting the real-world demands of this high-stakes field.
@@
By referring to the literature review below, several essential study gaps in ML-based IDS research have been noted. Still, there is enough opportunity to improve the accuracy of the intrusion detection model. Moreover, IDS users still struggle to detect known unencrypted attacks because IDS has not yet achieved 100% accuracy. One significant issue with IDS is that they regularly alert the auditor to false positives. However, an IDS can be optimised to reduce the number of false positives and false negatives, which is a deficiency of publicly available VANET databases with attacks. No fast, inexpensive, and effective in-built privacy-preserving ML-based IDS is still visible. VANET is a type of wireless network where both stationary and mobile nodes can connect. Such a system has frequent disconnections, a dynamic topology, and significant node mobility. Each attribute may result in a security flaw that hazards the system. Due to the high mobility of nodes, VANET topology is dynamic and unpredictable; this and other circumstances, such as severe weather conditions, may cause VANET to have several disconnections throughout the system.
@@ 1.1 Contributions
The contributions of this paper are as follows:
@@
    This research proposes a complimentary ensemble approach that integrates Random Forest, CatBoost, and Logistic Regression, providing a suitable trade-off between accuracy, explainability, and runtime requirement for intrusion detection within SDN-based VANETs.
@@
    We demonstrate our approach against the Vehicular Reference Misbehaviour (VeReMi) dataset to illustrate the recognition of different attack classes within VANETs. Furthermore, the integration of XAI using SHAP analysis, provides an explanability of how individual models and their collective ensemble can inform decision-making, and in particular, for understanding the model reasoning behind remaining misclassification. We chose Random Forest and CatBoost because they are particularly good at handling high-dimensional and categorical data. This capability is crucial for managing the complexity and variability of data found in VANETs.
@@ 2 Background
VANET, a type of wireless network connecting both stationary and mobile nodes, is characterised by frequent disconnections, a dynamic topology, and high mobility, which can lead to potential security flaws. Compared to Mobile Ad Hoc Networks (MANET), VANETs exhibit considerably greater mobility and more stringent topology restrictions. Vehicular Ad Hoc Networks (VANETs) enable applications such as real-time traffic monitoring, blind crossing prevention, dynamic route planning, collision prevention, and internet access for vehicles. According to the World Health Organization (WHO), 1.3 million people die in road crashes annually, with 20 to 50 million suffering non-fatal injuries (WHO (2023)).
The new era of automobiles and road traffic systems will depend on wireless networks and machine learning. As stated previously, many lives have been lost on highways. Consequently, when road traffic increases, road safety must also grow. We need a mechanism by which vehicles can be made sufficiently intelligent to manage their process for road safety. This notion constituted the foundation for VANETs. SDN-based VANETs follow the below architecture (Ku et al. (2014)).
2.1 Communications And Standards
The USDOT (The United States Department of Transportation) has designated Dedicated Short-Range Communication as the preferred inter-vehicle communication medium. The communication infrastructure for linked automobiles is provided via WAVE protocols. IEEE 802.11p protocol and IEEE 1609.x protocols define WAVE. Their activities conform to the internet protocol stack. Physical and data connection layers are supported by IEEE 802.11p. The tiers of operation for IEEE 1609.x are network, transport, and application. This protocol focuses on the physical and Media Access Control (MAC) levels to provide low latency and good dependability across short radio communication lines. It functions at both the physical and MAC levels. It offers a direct wireless interface and mapping between MAC and physical data. In addition, it provides media access rules and aids in multi-channel operations at the MAC layer (Chekkouri et al. (2015)).
@@ 2.1 Communications And Standards
IEEE 802.11p enables security via authentication and encryption of private communications. All cars are expected to be registered by a Motor Vehicle registration unit. This would assign MAC IDs to automobiles. Alternatively, Vehicle Identification Numbers (VINs) can serve as MAC addresses for automobiles. Vehicles will obtain IP addresses upon network connection. Vehicle identification will be based on their MAC addresses, IP addresses, or publicly provided IDs (Zheng and Wu (2016)).
@@
As highlighted in (Chen et al. (2009)), IEEE 1609.4p support for multi-channel operations is achieved by time division so radios can switch between different channel intervals. Control Channel (CCH) and Service Channel (SCH) are the two channels of this IEEE standard. Continuous safety messages are sent via CCH, whilst other applications are routed via SCH. IEEE 1609.3p defines the network and transport layer services. Here, IP addressing and routing occur. IEEE 1609.2p is the protocol that specifies secure message exchanges and encryption. IEEE 1609.1p is a protocol specifying alert and probe vehicle message forms. Currently, a modification to IEEE 1609.x is in the works. The following enhancements have been made. Draft IEEE 1609.5 to handle vehicle-to-vehicle and vehicle-to-infrastructure connectivity. Draft IEEE 1609.6 to provide remote management between Onboard Units and Road-Side-Units. IEEE Standard 1609.11 for safe electronic transactions.IEEE Standard 1609.12 defines the allocation of value identifiers in WAVE.
2.2 SDN strategies to address VANET security
@@ 3 Related Works
As a part of the research, different articles have been reviewed to explain how earlier researchers have ended work on an Intrusion Detection System for cyber-attack detection in SDN-based VANET. This Literature review study employs the methods stated by Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA). The methodology covers eight main steps: User needs assessment, define the Research Questions, Search String, Selection of Source, Select Relevant Paper, Screening Findings, Extract Data, and finally, synthesise the collected data (Kitchenham and Charters (2007)).
IDS and Malware detection systems can still not detect unknown, Encrypted, and Inference attacks with 100% accuracy. According to Kaspersky, most Malware detection products do not even reach the required 90% security threshold. A significant player in the IDS market, CISCO, claimed that Rapid detection—those made in less than five minutes—and long-term learning capacities displayed excellent performance in their IDS. Cisco Encrypted Traffic Analytics, or IDS, discovered nearly two-thirds of all known malicious flows in less than five minutes. Moreover, Cisco declared this IDS demonstrated better identification results even with low volumes of 0 to 20 flows, whereas 2000 plus flows enabled 100% detection for known attacks. Nevertheless, this solution has a limitation with a small attack size and taking three hours to gain 100% accuracy. However, the Cisco Adaptive Security Device Manager (ASDM) to manage Cisco Secure Firewall ASA (with IDS/IPS) has been impacted by Common Vulnerabilities and Exposures CVE-2022-20829, CVE-2021-1585, and CVE-2022-20828. Researcher Jake Baines, who disclosed them to Cisco in February and March of 2022, added that many users were undoubtedly not updating their Cisco firewalls properly (CISCO (2022) and Scroxton (2022)).
3.1.1 VeReMi
According to van der Heijden et al. (2018), the VeReMi dataset was created to test the effectiveness of VANET misbehaviour detection systems, based on the open-source VEINS vehicular simulation framework. This dataset can be efficiently employed to test the performance of misbehaviour detection approaches implemented on a city-wide scale while utilising fewer computing resources than standard (frequent) VEINS simulations. The dataset comprises On-Board Unit (OBU) message logs, including timestamps, sender IDs, positions, speeds, and reception times. VeReMi also includes a verification file, which provides GUI parameters on the validity of the data of each message upon which the verification confirms whether it is a “normal” or “attack” case. The dataset spreads across three attacker densities, five distinctive attack types, and three levels of density; it also includes various scenarios at different times of the day, including how traffic varies across the time of the day with the vehicle density range from 35 vehicles in low-dense areas to 519 vehicles during maximum congestion; and diverse types of attacks where the attacks in this paper are attacks against the position expressed in many ways from the position remaining still to noising out the position. While versatile, the VeReMi dataset is restricted in terms of both position-falsifying attacks and all types of network threats and advanced malicious strategies. Being a synthesised dataset, it lacks real-life cases but can be developed further by researchers.
@@ 3.1.1 VeReMi
The dataset is publicly available for reproducibility; thus, the research results made from the dataset are reliable and verifiable. VeReMi has 562,467 entries, out of which 437,429 entries are benign and 125,038 are attacking entries. The dataset comprehensively covers a city-level context of various types of roads and traffic behaviours, thus making it a reliable framework for benchmarking and testing the performance of misbehaviour detection approaches in VANETs (van der Heijden et al. (2018)). Table 1 shows the data classifications and the number of instances that occur within the dataset.
@@ 3.1.2 Attack detection techniques
The work by Singh et al. (2018) discusses the ML-based tactics to detect attacks in SDN-based vehicular networks based on machine learning techniques. The research has a particular emphasis on vehicle-to-infrastructure (V2I) links. It analyses the current environment to establish the ML strategy that will most successfully detect threats posed by distributed denial of service (DDoS) attacks (Short-lived Transmission Control Protocol Flooding, User Datagram Protocol Flooding). To accomplish this, it collects seven different elements of actual data from TCP (Transmission Control Protocol) and UDP (User Datagram Protocol) flows both when DDoS attacks them and when they are not. The network is educated by employing several different supervised learning algorithms depending on the data acquired, and it is discovered that the gradient boost classifier produces the most satisfactory outcomes. In addition, the random forest, decision tree, and linear support vector machine (linear-SVM) approach all produce results that are close to being the best. On the other hand, the limited training data could be to blame for the weak performance of Neural Networks (NN). Accurate, measurable data was not found in the write-up or the chart.
@@
According to Aneja et al. (2018), they presents an AI-based system to identify a Route Request (RREQ) flooding attack. Researchers used a genetic algorithm (GA) to optimise feature subset selection and an Artificial Neural Network (ANN) for classification. The model scored 99% accuracy from Network Simulator 2 (NS2) 1000 entries simulated dataset.
@@
In their study, Yu et al. (2018) analyse DDoS assaults in an SDN-based vehicular network by employing TCP flood, UDP flood, or ICMP (internet control message protocol) flood. In this research paper, the authors present a scheme for detecting Distributed Denial-of-Service (DDoS) attacks in SDVN atmospheres using three separate models. A “trigger detection model” for inbound packets and another method is the “flow table feature-based detection model”. Both methods use the OpenFlow protocol features and the detection of a threat model based on support vector machine classification. All of these models are used in conjunction with one another. They synthesised and simulated network traffic to accommodate various DDoS attacks via Scapy and hping3, and the findings suggested an efficiency of over 97%. In the article, ML-Based Intrusion Detection Method for VANETs.
@@
In their research, Zeng et al. (2018) developed an ML-based Senior2Local IDS with propagation to identify potential threats in VANETs both Globally and Locally. They used a fuzzification ANN technique to recognise malicious attacks globally and protect Cluster Heads in Roadside Units (RSUs). After that, a game theory trust system or RSU is developed separately. They employed a simple and straightforward Support Vector Machine (SVM) to detect attacks locally. The findings demonstrated that the method was more reliable and resilient than earlier ML-based tactics.
@@
As outlined by Gad et al. (2021), the researchers conducted the experiment on a ToN-IoT dataset and tested various ML approaches in binary and multiclass classification problems in VANET to detect intrusion. On the other hand, LR is used for solving binary classification problems, while multiclass issues are addressed with a one-vs-rest approach.
@@
3.1.3 Methods yet to reach optimum accuracy
As reported by Türkoğlu et al. (2022), with the help of feature selection and hyperparameter optimisation, the DT classifier attained an accuracy score of 99.35%, which was the best possible result. The excellent results achieved with the methodology developed proved the considerable significance of feature selection and optimisation techniques. These findings were demonstrated by the outstanding results obtained with the method.
In the findings of Kim et al. (2017), a collaborative security threat detection method is demonstrated within the software-defined vehicular cloud architecture. This method highlights the importance of collaborative approaches in enhancing security within vehicular networks. A traditional vehicular environment has few exceptional features, such as high flexibility, dynamic networking, and rapid connectivity. His proposed model said a Speed Dependent Volume Controller (SDVC) had collected the flow information from the vehicular cloud (VC) to train multiclass SVM. In the next step, the SVM classifier identifies the bad actors among the network’s nodes. However, performance is still under 90% for Accuracy score, Precision, and Recall.